[Ncep.nhc.nco_contacts] TSB Morning Rounds - Thu Feb 11, 2016

Mark DeMaria - NOAA Federal mark.demaria at noaa.gov
Thu Feb 11 12:42:25 UTC 2016


Here is a summary of today's computer operations.

Mark

-----

TAFB:

TSB is continuing to work with NCO on getting ftp gate access from the
debout standing desk N-AWIPS system.

NCO is finished configuring the hardware replacement of the derecho N-AWIPS
system at the other standing desk. Since Martin is off tomorrow and Monday
is a holiday, we are going to wait until later next week for the swap.

There was an ASCAT data outage yesterday due to problems with a NESDIS
server. That issue has been resolved.

HSU:

No new issues.

Other:

NCO confirmed that the short internet outage around noon on Tuesday was
caused by a firewall in College Park that needed an emergency reboot. The
problem was fixed before we had a chance to report it to them.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160211/f74249bb/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list